A car starts from rest and accelerates uniformly over a time of 5.21 seconds for a distance of 110 m. Determine the acceleration of the car.



Grade:11

2 Answers

Ambuj Bansal
10 years ago

s=ut+at^2/2

110=0+a27.14/2

220/27.14=a

a=21.4m/s^2

 

Smile

SUBHADEEP ROY
10 years ago

8.1048 m/s2